how to find rational numbers between two rational numbers

Grade:7

1 Answers

Harshit Panwar
6 years ago
Add those two numbers and divide them by 2. Or find mean of both the numbers.
Example :-
3 and 9
Mean of 3 and 9
(3+9)/2=12/2=6
First number is 6.
Second number is (3+6)/2=4.5
Similarly third number (6+9)/2=7.5
Similarly forth number (3+4.5)/2=3.75
 
Hope it helps you.
Approve if it helps you.